# Kinetic Interacting Particle Langevin Monte Carlo

### v2

@@ B.5 Proof of Theorem 2
$\tilde{C}=\sqrt{3}(\sqrt{NL}\lor\sqrt{NL}^{-1}),\qquad\tilde{K}=NL\left(1+e^{NL\tilde{\eta}^{2}}\left(\frac{\tilde{\eta}}{6}+\frac{\tilde{\eta}^{2}NL}{24}\right)\right)\left(1+\frac{\tilde{\eta}NL}{2\sqrt{\tilde{\mu}}}\right).$
Again, we can recover the behaviour of KIPLMC2 when we set $\tilde{\eta}=\eta/\sqrt{N}$ and $\tilde{\gamma}=\sqrt{N}\gamma$. It is quite easy to see that when we set these parameters and consider the $\theta$-marginal, the $n$th step of this rescaled scheme will now correspond to the $n$th step of the KIPLMC2, provided the initial distribution of both is the same. Combining this with the concentration result from Prop. 2, we obtain the desired result.
Appendix C Results
C.1 Area Between Curve Calculation (ABC)
@@
To compare the behaviour of the different algorithms, we seek to identify a measure that quantifies how accurately and quickly each algorithm converges to the true solution. To do this we employ the ABC to do this, as in [49]. Consider $c^{k}:\mathbb{N}\to\mathbb{R}$ and $c^{m}:\mathbb{N}\to\mathbb{R}$ to be the accuracy curves of KIPLMC2 and MPGDnc respectively, i.e. for $\{\theta_{n}\}_{n=1}^{\infty}$ generated by the KIPLMC2 algorithm, $c^{k}(n)=\|\theta_{n}-\bar{\theta}^{\star}\|$. The ABC we will compute is,
@@
where $M$ is the total number of time-steps and $w$ acts as a normalising, weighting function. We set $w(i)=\frac{2}{M(M+1)}\frac{i}{M}$ for our computations, as it matches the weighting used in [47]. Hence, the ABC computes a signed, weighted area between the error curves of MPGD and KIPLMC2. In particular, when $c^{m}$ is dominated by $c^{k}$, the ABC is negative, whilst positive in the converse case.
C.2 Variance from (c) Fig. 3 in tabular form
Table 2: We compare the performance of the different algorithms over increasing step-sizes. Observe, despite the better behaviour at lower step-size $\eta$, the MPGDnc algorithm is the least stable. We compute the sample variance of the algorithms over 20 Monte Carlo simulations, using the last 500 steps of each simulation.
Sample Variance
Algorithms $\eta$
0.01 0.025 0.05 0.075 0.1
MPGDnc $\bm{2.30\times 10^{-2}}$ $\bm{2.77\times 10^{-2}}$ 1.75×$\times 10^{9}$ 2.34×$\times 10^{50}$ 7.78×$\times 10^{110}$
KIPLMC1 1.91×$\times 10^{-1}$ 1.30×$\times 10^{-1}$ 1.90×$\times 10^{-1}$ 8.52×$\times 10^{-1}$ 6.06×$\times 10^{17}$
KIPLMC2 2.63×$\times 10^{-2}$ 3.05×$\times 10^{-2}$ $\bm{3.37\times 10^{-2}}$ $\bm{3.03\times 10^{-2}}$ $\bm{3.36\times 10^{-2}}$
C.3 Error Metrics for the BNN
@@
Indeed, by [47], assuming that the data generating process produces independent samples, then this metric approximates the difference of the Kullback-Leibler divergence and the entropy of the data generating process. Indeed, the LPPD approximates
@@
where $p(l,f)$ is density from which the data is independently sampled. Hence, the larger the LPPD, the smaller the Kullback-Leibler divergence between our estimate and the true distribution.

### v3

@@ C.5 Proof of Theorem 2
${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq\tilde{C}_{1}\left(1-\frac{\eta\mu}{3\gamma}\right)^{n/2}+\eta\tilde{C}_{2}+\sqrt{\frac{\tilde{C}_{3}}{N}},$
with $C_{3}=d_{\theta}/\mu$.
Appendix D Results
D.1 Multi-well Experiment
We include the following multi-well experiment to highlight the advantage of the noised $\theta$ dynamics, the core difference between the proposed methods and MPGD [37]. To do this, we consider a synthetic example with non-convex posterior. Consider the joint probability density over $\mathbb{R}$,
$p_{\theta}(x,y)$ $=\mathcal{N}(y;x,\sigma^{2}_{y})p_{\theta}(x),$
$p_{\theta}(x)$ $=\frac{1}{2}\mathcal{N}(x;-\theta,\sigma^{2}_{x})+\frac{1}{2}\mathcal{N}(x;\theta,\sigma_{x}^{2}).$
One can note that given small enough choices of $\sigma_{x}$ and $\sigma_{y}$, as well as, large enough $\bar{\theta}_{\star}$, this density has a multi-modal $\theta$-marginal. Indeed, let us observe that the parameter posterior is given as,
$p_{\theta}(y)=\frac{1}{2}\mathcal{N}(y;\theta,\sigma_{x}^{2}+\sigma_{y}^{2})+\frac{1}{2}\mathcal{N}(y;-\theta,\sigma_{x}^{2}+\sigma_{y}^{2}).$
In this case, we seek to show that whilst the MPGD algorithm gets stuck in one of the local minima, the KIPLMC1 algorithm escapes these under the same hyperparameters. For the experiment we generate 1 data-points at $\theta=1.5$, from which we obtain true $\bar{\theta}_{\star}=\pm 1.5$, and we fix $\sigma_{x}^{2}=1$ and $\sigma_{y}^{2}=0.25$. In Fig. 5, we show the trace plots of $\theta$, together with the associated empirical density plots, compared to $\bar{\theta}_{\star}$ and the marginal likelihood $p_{\theta}(y)$ respectively. For sake of comparison, we also compare the dynamics of KIPLMC1 and KIPLMC2 on the marginal dynamics, i.e. we consider KIPLD to be given simply as
$\mathrm{d}\bm{\theta}_{t}$ $={\mathbf{V}}_{t}^{\theta}\mathrm{d}t,$
$\mathrm{d}{\mathbf{V}}^{\theta}_{t}$ $=-\gamma{\mathbf{V}}_{t}^{\theta}-\nabla_{\theta}\kappa(\theta)\mathrm{d}t+\sqrt{\frac{2\gamma}{N}}\mathrm{d}{\mathbf{B}}_{t},$
for a $\mathbb{R}^{d_{\theta}}$ Brownian motion ${\mathbf{B}}_{t}$ and we recall from Proposition 1 that $\kappa(\theta)=-\log p_{\theta}(y)$. In Fig. 5 we observe that all methods converge close to the true solution, but only the “noised” algorithms KIPLMC1, KIPLMC2 and IPLA successfully have support on both true solutions reliably.
Figure 6: Runtime Comparison. We compare the MSE for the synthetic logistic regression experiment performed in Sec. 5.1 with the same hyperparameters $d_{x}=d_{\theta}=3$, $d_{y}=100$, $\gamma=2.2$ and $N=100$. The results shown are the average of 100 Monte Carlo simulations.
D.2 Computational Cost Comparison
We further perform an additional experiment on the Sec. 5.1 example, to compare MSE against computational cost. Indeed, this allows for more thorough comparisons with the methods, as KIPLMC1 generates correlated noise, KIPLMC2 performs a gradient correction and MPGD does both. Further all these methods require twice as much space in memory, when compared to PGD and IPLA, due to the presence of the momentum variables.
In Fig. 6, we observe the faster initial convergence rate of the momentum based methods, despite their greater computational cost. Further, we observe that the increased cost of noising the $\theta$-dynamics in our proposed methods, is negligible when compared to a second gradient evaluation per iteration, required by MPGD and KIPLMC2.
D.3 Area Between Curve Calculation (ABC)
@@
To compare the behaviour of the different algorithms, we employ a measure that quantifies how accurately and quickly each algorithm converges to the true solution, namely the ABC, as in Lim et al. [37]. Consider $c^{k}:\mathbb{N}\to\mathbb{R}$ and $c^{m}:\mathbb{N}\to\mathbb{R}$ to be the accuracy curves of KIPLMC2 and IPLA respectively, i.e. for $\{\theta_{n}\}_{n=1}^{\infty}$ generated by the KIPLMC2 algorithm, $c^{k}(n)=\|\theta_{n}-\bar{\theta}^{\star}\|$. The ABC is given as,
@@
where $M$ is the total number of time-steps and $w$ acts as a weighting function $w(i)=2i/(M(M+1))$, matching Kuntz et al. [35]. Hence, the ABC computes a signed, weighted area between the error curves of IPLA and KIPLMC2. In particular, when $c^{m}$ is dominated by $c^{k}$, the ABC is negative, whilst positive in the converse case.
D.4 Error Metrics for the BNN
@@
Indeed, by Kuntz et al. [35], assuming that the data generating process produces independent samples, then this metric approximates the difference of the Kullback-Leibler divergence and the entropy of the data generating process. Indeed, the LPPD approximates
@@
where $p(l,f)$ is the density from which the data is independently sampled. Hence, the larger the LPPD, the smaller the Kullback-Leibler divergence between our estimate and the true distribution.